The wavelength of light visible to the human
eye is on the order of 5 × 10−7 m.
Find the frequency of the lightwave if the
speed of light in air is 2.99792 × 108 m/s.
Answer in units of s−1.
1
Expert's answer
2020-02-17T09:23:35-0500

The frequency of the electromagnetic wave


"f=\frac{c}{\lambda}""=\frac{2.99792\times 10^8\:\rm m/s}{5\times 10^{-7}\:\rm m}=6\times 10^{14} \:\rm s^{-1}."